OpenStudy (anonymous):

3x-2y=5 -6x+4y=7 Solve the linear equation algebraically!

OpenStudy (anonymous):



OpenStudy (anonymous):

can anyone answer please?

OpenStudy (anonymous):

try using substitution or elimination

OpenStudy (anonymous):

i know that. i dont understand HOW to use it though.

OpenStudy (anonymous):

okay take one of the equations, anyone and make it equal x or y ur choice. for ex: 3x-2y=5 +2y =2y 3x=5+2y /3 /3 x=5/3+2/3y Now take the other equation and subsitute the x in that equation for 5/3+2/3y and solve.

OpenStudy (anonymous):

3x-2y=5..................(1) -6x+4y=7............(2) y=(-5+3x)/2 [ from equation 1] sub y in (2) -6x+4[(-5+3x)/2=7 0=17this means x is contained in empty set hence the system of linear equations is inconsistence(no solution)